OpenStudy (anonymous):

Write the expression sin(tan^-1)2x)

OpenStudy (dumbcow):

\[\theta = \tan^{-1} 2x\] \[\tan \theta = 2x\] we want to find sin theta using triangles .... tan = opp/adj = 2x/1 ** sorry its not letting me use the draw feature** hyp = sqrt(1+4x^2) sin theta = 2x/sqrt(1+4x^2)

OpenStudy (anonymous):

thnk you!!

OpenStudy (dumbcow):

yw :)